# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

### v1

@@ 3.2 Application Scenarios
High Generalizability Tool using • General tool using • Multimodal tool using  [127, 128, 129]  [130, 131]
@@
Dialugue Agents • Conducting persuasive dialogues • Dialogue task optimization  [132, 133, 134, 135]  [136, 137, 138, 139]  [140, 141, 142, 143]  [144, 145]
@@
Multimodal Perception • Multimodal tool using • Instructions acceptance, image capture,  [146, 127, 128, 147]
@@
Multi-Agent Collaboration • Multi-agent social interaction modeling • Collaborative task division and quality  [148, 149, 150, 151]  [152, 153, 154, 155]  [156, 157, 158]
@@
Visual Tasks Processing • Reasoning covered object in images • Image Inpainting  [159, 160, 161, 162]
@@
Natural Language Processing • Keyphrase generation • Long text reading comprehension  [163, 164, 165]  [166, 167, 168]
@@
Data Generation and Analysis • Transporing the knowledge of the isolated agents  [169]
Reasoning, Planning and Decision Optimization • Autonomous information retrieval in visual • Robotic language-guided task planning  [170, 171, 172, 173]  [174, 175, 176, 177]  [178, 179, 180, 181]  [182, 183, 184, 185]  [186, 187, 188, 189]  [190, 191, 192]
@@
Limited Generalizability Function-driven Code Generation • Multi-agent code generation framework for • Code generation and environment interaction  [193, 194, 195, 196]
@@
Social Simulation • User behavior simulation • Human social interaction simulation  [197, 198, 199, 200]  [58, 201, 202, 203]  [204, 205, 206, 207]  [208, 209, 210]
@@
Graphical User Interface Agents • Multi-turn conversational web agents • Personalized web agents  [211, 212, 213, 214]  [215, 216, 217, 218]  [219, 220, 221, 222]  [223, 224, 225, 226]  [227, 228, 229]
@@
Model Analysis, Evaluation and Improvement • Dialogue task optimization • Multi-agent evaluation optimization  [230, 62, 231]  [232, 233, 234]
@@
Scenario-driven Finance • Multimodal foundation agent for financial trading • Financial sentiment analysis  [235, 236, 237, 238]
@@
Science • Automating data science tasks • Automating scientific discovery  [239, 61, 240]  [241, 242]
@@
Healthcare • Medical reasoning and diagnosis • Electronic health record modeling and  [243, 244, 245, 246]  [247, 248, 249]
@@
Game • Fighting game agents training • Intelligence evaluation in murder mystery  [250, 251, 252]  [253, 254]
@@
Manufacturing • Dynamic scheduling in flexible manufacturing  [254]
@@
Robotics • LLM-based embodied agent open environment • Embodied instruction following  [255, 256, 257, 258]
@@
Urban Computing • Addressing standardization tasks supported • Urban knowledge graph construction  [259, 260, 261]
@@
Social Media • Detecting harmful content on social media • Logical and causal fact-checking  [262, 263]
Autonomous Driving • Real-time motion generation for autonomous driving • Traffic signal control  [264, 265]
Literary Creation • Director-actor coordinate agent framework for • Narrative generation through multi-step  [266, 267]
@@
Given the poweful knowledge acquisition, instruction understanding, task planning and logical reasoning capabilities of multi-agent systems, they have been widely used worldwide. From small and medium-sized organizations to multinational corporations, from higher education institutions to research institutes, from grassroots medical facilities to large diagnostic hospitals, and government agencies at all levels are actively promoting the application of AI agent technology. In terms of application scenarios, the penetration of AI agent technology is extremely broad, covering multidisciplinary fields such as healthcare, education, trade and economics and law. The problems that multi-agent systems are capable of addressing ranges from comprehending a single piece of literature to simulating global-scale social events, spanning across micro, meso and macro levels. In contrast to categorizing applications based on disciplinary domains [14] or task types [112], we sort and categorize application scenarios according to the degree of generalizability of multi-agent systems and build upon this to explore the associated issues of value alignment.
Generalizability of multi-agent systems refers to the extent to which they can be applied across diverse scenarios or meet the needs of various organizations. A high degree of generalizability is based on the foundational capabilities of AI agents such as natural language processing, reasoning, planning and autonomous decision-making. The universal application includes tool using, dialog agents, multimodal perception, multi-agent collaboration, visual tasks proingcess, natual language processing, data generation and analysis, reasoning, planning and decision optimization and on, see Table II.
In terms of tool using, AI agents are capable of recognizing implicit needs for tool invocation and leveraging a wide range of real-world tools involving perception, manipulation, logical reasoning, and creativity. These tools enable AI agents to process multimodal inputs—such as images, tables, and code—to accomplish complex tasks [129, 130]. Taking dialogue agents as an example, A.V. Savchenko and L.V. Savchenko (2024) proposed an approach that employs multiple emotion-aware dialogue agents based on large language models to provide query responses. These AI agents integrate speech and facial expression recognition to infer the user’s emotional state and generate appropriate responses [134]. Furthermore, dialogue agents have also been applied in persuasive interactions for social good, particularly in scenarios involving user resistance, where they are able to generate more polite and convincing responses [132]. In the domains of reasoning, planning, and decision-making optimization, autonomous agents can guide large language models to perform zero-shot reasoning, thereby improving their performance in generation, classification, and inference tasks  citeDBLP:conf/icml/CrispinoMZS024. Multi-agent debate mechanisms have been shown to enhance accuracy and factual consistency in mathematical and strategic reasoning while reducing hallucinated outputs [175]. Additionally, structuring multi-agent interactions as graph representations enables knowledge distillation for smaller models, improving their performance and generalization in commonsense and mathematical reasoning tasks [176].
Higher degree of generalizability of indicates that the agent system can be broadly applied across various scenarios, not only addressing domain-specific problems but also meeting diverse task requirements across multiple industries and organizations. For example, multi-agent systems can be applied not only in natural sciences but also in social science domain such as economics and management, suitable for addressing global issues that span across nations and cultures.
Lower generalizability means that multi-agent systems can not be applied across industries and domains. Specialized agent applications can be broadly categorized into two types: function-driven and domain-driven, see Table II.
@@
Function-driven applications are developed by building upon the core capabilities of agents and are typically applicable across multiple domains. Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model analysis, evaluation and improvement, as in Table II. For instance, in social simulation, Wang L. et al. (2023) proposed a large language model-based agent framework that simulates user behavior in sandbox environments to examine social phenomena such as echo chambers and herding effects [197]. In GUI navigation, agents can handle open-domain web tasks by extracting key information from complex HTML pages, understanding web structures, and performing multi-step interactions, thus addressing challenges like interaction diversity and domain knowledge scarcity [225, 227]. In model analysis, LLM-based agents have been employed to assess alignment performance, such as in Ali-agent, which simulates test scenarios and iteratively explores long-tail risks to evaluate value alignment [231]. MAIA, a multimodal explanatory agent, automates model behavior interpretation using neural networks, enhancing model transparency [62].
@@
Scenario-driven applications refer to agent systems deployed within specific sectors, including finance, science, healthcare, game, manufacturing, and autonomous driving, as in Table II. In urban computing, the UrbanKGent framework enables automated construction of urban knowledge graphs, significantly reducing development costs [261]. In scientific research, agents assist with academic literature retrieval and citation recommendation [61], organic structure design [241], and automated data visualization. In healthcare, multi-agent systems support medical reasoning, gamified mental health assessments, and experimental design in genomics [243, 247]. In finance, agents can simulate macroeconomic dynamics [236], analyze market trends using multimodal data [235], and support trading and portfolio management [237]. Furthermore, in the social media domain, agent reasoning and coordination capabilities contribute to tasks such as harmful content detection and high-confidence fact verification [262, 263].
@@
The low level of generalizability indicates that the agent system is used in specific countries, industries, or even organizations and enterprises. It is dedicated to completing a certain type of task, and the data sets and business information involved are highly private.
3.3 Value Issues in Application
The value alignment problems involved in the application scenarios of LLM-based agent systems mainly include two aspects: the correspondence between the generalizability of the system and the multi-level values, and the contradictions encountered when aligning values.
The degree of generalizability of LLM-based agent systems corresponds to the multi-level of values. As the degree gradually decreases, the granularity of the value norms that need to be met gradually increases. A higher degree of generalizability indicates that the applicable scenarios for LLM-based agent systems are more diverse. Therefore, it is unnecessary for the initial design of systems to align with the values of a specific group–such as a country, industry, or culture–to ensure flexible applicability across national, ethnic, and industrial contexts. The development and governance of highly generalizable multi-agent systems should be constrained by universal values of macro level. Multi-agent systems with limited generalizability typically confined to specific countries or industries. Such systems often exhibit design features characteristic of particular countries, industries, or cultures. Therefore, the development and governance of specialized systems must address diverse national policies, industry standards, and cultural preferences, while ensuring alignment with universal values. When intelligent systems are deployed in strictly private contexts, they are typically confined to specific organizations and designed to address specialized business or personal needs. These applications are highly specialized and exhibit minimal generalizability. Under such conditions, it is essential not only to align with value principles of macro and mecro level, but also to comply with the specific value frameworks of the organizations or businesses involved. It can be seen that multi-agent systems with high generalizability tend to be highly adaptable, and their value alignment primarily centers on adherence to universal principles of macro level. As application specialization increases, the requirements for value alignment become progressively more complex, necessitating adherence to value norms across multiple levels simultaneously. This classification aligns with the top-down, increasingly granular hierarchy of value principles. As a result, the value norms that multi-agent systems must adhere to vary according to their degree of generalizability, offering a structured basis for governance and oversight in their design and deployment.
The value alignment problem may involve two tpyes of tensions: the plurality of value principles at the same level, and potential conflicts between values across different levels. On one hand, a key challenge is that value principles at the same level differ across countries, regions, industries, organizations, and other contextual factors. Therefore, when multi-agent systems are applied in diverse contexts, their value alignment objectives must be adapted accordingly. In the healthcare sector, Huawei’s intelligent agent EIHealth serves as a professional AI development platform for genomic analysis, drug discovery, and clinical research. Built on Huawei Cloud’s strengths in AI and big data, EIHealth features an open and extensible architecture that enables end-to-end AI empowerment. Ensuring alignment with the value principles and regulatory norms of various countries, regions, and organizations is crucial for its deployment. To address this, Huawei Cloud has issued a series of white papers tailored to the privacy and regulatory requirements of different jurisdictions and industries. For example, it provides compliance guidelines aligned with the Personal Data Privacy Ordinance (PDPO) in the Hong Kong SAR, as well as guidelines addressing the financial regulatory frameworks of the EU, including specific provisions for Ireland, Spain, and Hungary. These efforts ensure that the EIHealth platform adheres to all applicable national and regional security regulations, international cybersecurity and cloud security standards, and incorporates industry best practices—ultimately meeting the diverse security needs of cloud service users.
On the other hand, as the tasks assigned to an intelligent system become more fine-grained, the constraints on achieving alignment goals also increase. The means of maximizing task utility may conflict with established value norms, leading to tensions across different levels of values. For instance, the promotion of social cohesion at the macro level may contradict individualistic tendencies at the meso level, where individuals prioritize self-interest. Similarly, national regulations that emphasize information transparency may be at odds with micro-level demands for organizational data privacy. These cross-level value conflicts pose significant challenges for the design and development of multi-agent systems. In the field of corporate recruitment, algorithms and artificial intelligence have been widely adopted. Recent studies have explored the design and development of intelligent agent systems for tasks such as résumé screening and candidate-job matching [268]. However, AI-based recruitment systems may be at odds with widely accepted ethical norms [104]. In particular, corporate hiring practices can involve gender discrimination or racial bias, which directly conflict with legal regulations and core societal values. This ethical dilemma was vividly illustrated by Amazon’s 2018 algorithmic bias incident. The company’s internally developed résumé screening system, trained on historical hiring data, systematically discriminated against female applicants, ultimately leading to the discontinuation of the system [269]. To prevent such issues, recruitment-oriented intelligent systems must be explicitly designed to reduce or eliminate gender and racial bias. Their value alignment should comply with national labor laws, equal employment regulations, women’s rights protections, and relevant industry standards. Only on the basis of legal compliance, equality, and privacy protection should these systems then pursue the optimization of recruitment outcomes.
In complex collective intelligence systems, multi-agent value alignment refers to the design of appropriate mechanisms and algorithms that enable multiple autonomous agents to identify and adhere to shared ethical or social norms while pursuing their individual task objectives, thereby achieving consistency, safety, and fairness at the system level. Unlike single-agent alignment, which focuses on the compliance of a model’s outputs, multi-agent collaborative alignment must address challenges such as value conflicts between agents, information asymmetry, and “free-rider” or competitive dilemmas in cooperative games. The goal is to use communication protocols, reward shaping, negotiation mechanisms, or adversarial training to ensure that agents coordinate with one another and ultimately converge on a common set of collective values.
4 Evaluation and Methods for Value Alignment
To ensure that LLM-based agents behave in ways consistent with human values and task goals, value alignment evaluation has become a key research focus. Evaluation methods often rely on alignment benchmarks to assess whether agents follow expected value principles. On the methodological side, alignment techniques can be applied at various stages, including during pretraining, post-training, and through reinforcement learning (e.g., reward modeling or fine-tuning with human feedback). These methods aim to mitigate misalignment risks and enhance agent trustworthiness across various applications.
4.1 Value Alignment Evaluation
4.1.1 Datasets for Value Alignment Evaluation
@@
Table III summarizes the existing open-source datasets for evaluating value alignment. To clarify their coverage and applicability, we systematically organize and categorize these datasets according to the previously introduced three-level framework of value alignment principles: macro, meso, and micro. This hierarchical classification enables a more nuanced understanding of how current resources support value alignment evaluation at different levels.
Table III: Statistics of Datasets for Value Alignment Evaluation.
AC: Automatic Construction, CC: Crowdsourced Construction, HC: Hybrid Construction, EDC: Expert-Driven Construction, IRED: Integration and Refinement of Existing Datasets, BJ: Binary Judgments, RT: Ranking Tasks, MCQ: Multiple-Choice Questions, OEQ: Open-Ended Questions, RSQ: Rating Scale Questions
Datasets Macro Level Meso Level Micro Level Construction Methodologies Question Formats Size
BBQ(Bias Benchmark for QA) [270] EDC MCQ 58.4k
BeaverTails [51] IRED OEQ 30.2k
BOLD [271] CC OEQ 23.6k
CBBQ(Chinese Bias Benchmark Dataset) [272] HC MCQ 106.5k
CDEval [75] AC MCQ 2.9k
CDial-Bias [273] CC OEQ 28k
CORGI-PM [274] IRED BJ, OEQ 32.9k
CrowS-Paris CC BJ 1.5k
CultureSPA [275] AC MCQ 13k
Cvalues [276] CC MCQ, OEQ 6.4k
DailyDilemmas [277] AC MCQ 2.7k
DecodingTrust [278] IRED MCQ 152.4k
DEFSurveySim [49] HC MCQ, RSQ 1.1k
EEC(Equity Evaluation Corpus) [279] EDC BJ 8.6k
ETHICS [280] CC MCQ, BJ, RT 134.4k
Flames [281] CC OEQ 1k
GlobalOpinionQA [282] IRED MCQ 2.5k
HofstedeCulturalDimensions [76] HC MCQ 0.2k
IndieValueCatalog [283] IRED MCQ 93.2k
KorNAT [284] AC MCQ 10k
LLMGlobe [77] HC RSQ, OEQ 37.6k
MFQ(Moral Foundations Questionnaire) [285] IRED MCQ 11k
Moral Beliefs [286] AC MCQ, BJ, OEQ 0.5k
Moral Integrity Corpus [287] AC RSQ, OEQ 38k
Moral Stories [288] CC MCQ 12k
MoralExceptQA [289] EDC RT 0.2k
NaVAB [290] AC MCQ, BJ 63.8k
Persona Bias [279] IRED MCQ, BJ 110.6k
PkuSafeRLHF [291] AC OEQ 476.4k
ProgressGym [292] HC RSQ, OEQ 1.4k
SafeSora [60] AC OEQ 14.7k
Scruples [293] EDC MCQ, BJ 657k
Social Bias Frames [294] CC RT, OEQ 150k
Social Chemistry 101 [295] CC BJ, OEQ 292k
StereoSet [296] CC RT, OEQ 4.2k
ToxiGen [297] AC OEQ 6.5k
UnQover [298] IRED BJ, OEQ 2713k
ValueNet [299] CC BJ, OEQ 21.3k
WikiGenderBias [300] IRED OEQ 45k
WinoBias [301] EDC MCQ 3.1k
WinoGender [302] EDC BJ, OEQ 0.7k
@@
Although existing datasets address multiple levels of value alignment, several limitations persist. At the macro level, many datasets—such as Moral Stories [288] and ETHICS [280]—are predominantly grounded in Western cultural contexts, lacking comprehensive representation of alternative value systems, such as those from Asian or African societies. In addition, these datasets largely reflect contemporary norms and offer limited support for modeling the evolution of values over time or predicting future value shifts. At the meso level, datasets such as NaVAB [290], KorNAT [284], and CultureSPA [275] provide valuable insights into national or cultural contexts. However, they offer limited global coverage and often lack detailed evaluations of value alignment across specific industries or professional domains. Micro-level datasets are even more scarce. Compared to macro- and meso-level resources, they remain significantly underdeveloped, which limits the ability to assess and align value preferences in multi-agent systems across diverse, fine-grained real-world scenarios.
4.1.2 Dataset Construction Methodologies

### v2

@@ Appendix A Application Scenarios of Agentic AI systems
High
Generalizability
Tool Using • General tool using • Multimodal tool using
 [231, 232, 209]
 [210, 233]
@@
Dialugue
Agents
• Conducting persuasive dialogues • Dialogue task optimization
 [212, 234, 211, 235]
 [236, 237, 238, 239]
 [240, 241, 242, 243]
 [244, 245]
@@
Multimodal
Perception
• Multimodal tool using • Instructions acceptance, image capture,
 [246, 231, 232, 247]
@@
Multi-Agent
Collaboration
• Multi-agent social interaction modeling • Collaborative task division and quality
 [248, 249, 250, 251]
 [252, 253, 254, 255]
 [256, 257, 258]
@@
Visual Tasks
Processing
• Reasoning covered object in images • Image Inpainting
 [259, 260, 261, 262]
@@
Natural Language
Processing
• Keyphrase generation • Long text reading comprehension
 [263, 264, 265]
 [266, 267, 268]
@@
Data Generation
and Analysis
• Transporing the knowledge of the isolated agents
 [269]
Reasoning, Planning
and Decision
Optimization
• Autonomous information retrieval in visual • Robotic language-guided task planning
 [270, 271, 213, 272]
 [273, 214, 215, 274]
 [275, 276, 277, 278]
 [279, 280, 281, 282]
 [283, 284, 285, 286]
 [287, 288, 289]
@@
Limited
Generalizability
Function-driven Code Generation • Multi-agent code generation framework for • Code generation and environment interaction
 [290, 291, 292, 293]
@@
Social Simulation • User behavior simulation • Human social interaction simulation
 [216, 294, 295, 296]
 [297, 298, 299, 300]
 [301, 302, 303, 304]
 [305, 306, 307]
@@
Graphical User
Interface Agents
• Multi-turn conversational web agents • Personalized web agents
 [308, 309, 310, 311]
 [312, 313, 314, 315]
 [316, 317, 318, 319]
 [320, 321, 217, 322]
 [218, 323, 324]
@@
Model Analysis,
Evaluation and
Improvement
• Dialogue task optimization • Multi-agent evaluation optimization
 [325, 220, 219]
 [326, 327, 328]
@@
Scenario-driven Finance • Multimodal foundation agent for financial trading • Financial sentiment analysis
 [227, 226, 228, 329]
@@
Science • Automating data science tasks • Automating scientific discovery
 [330, 222, 331]
 [223, 332]
@@
Healthcare • Medical reasoning and diagnosis • Electronic health record modeling and
 [224, 333, 334, 335]
 [225, 336, 337]
@@
Game • Fighting game agents training • Intelligence evaluation in murder mystery
 [338, 339, 340]
 [341, 342]
@@
Manufacturing • Dynamic scheduling in flexible manufacturing
 [342]
@@
Robotics • LLM-based embodied agent open environment • Embodied instruction following
 [343, 344, 345, 346]
@@
Urban
Computing
• Addressing standardization tasks supported • Urban knowledge graph construction
 [347, 348, 221]
@@
Social Media • Detecting harmful content on social media • Logical and causal fact-checking
 [229, 230]
Autonomous
Driving
• Real-time motion generation for autonomous driving • Traffic signal control
 [349, 350]
Literary Creation • Director-actor coordinate agent framework for • Narrative generation through multi-step
 [351, 352]
@@
Specialized applications can be broadly categorized into two types: function-driven and domain-driven. Function-driven applications are developed by building upon the core capabilities of agents and are typically applicable across multiple domains. Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model analysis, evaluation, and improvement. For instance, in social simulation, Wang L. et al. (2023) proposed a large language model-based agent framework that simulates user behavior in sandbox environments to examine social phenomena such as echo chambers and herding effects [216] In GUI navigation, agents can handle open-domain web tasks by extracting key information from complex HTML pages, understanding web structures, and performing multi-step interactions, thus addressing challenges like interaction diversity and domain knowledge scarcity [217, 218]. In model analysis, LLM-based agents have been employed to assess alignment performance, such as in Ali-agent, which simulates test scenarios and iteratively explores long-tail risks to evaluate value alignment [219]. MAIA, a multimodal explanatory agent, automates model behavior interpretation using neural networks, enhancing model transparency [220].
@@
Scenario-driven applications refer to agentic AI systems deployed within specific sectors, including finance, science, healthcare, game, manufacturing, and autonomous driving. In urban computing, the UrbanKGent framework enables automated construction of urban knowledge graphs, significantly reducing development costs [221]. In scientific research, agents assist with academic literature retrieval and citation recommendation [222], organic structure design [223], and automated data visualization. In healthcare, multi-agent systems support medical reasoning, gamified mental health assessments, and experimental design in genomics [224, 225]. In finance, agents can simulate macroeconomic dynamics [226], analyze market trends using multimodal data [227], and support trading and portfolio management [228]. Furthermore, in the social media domain, agent reasoning and coordination capabilities contribute to tasks such as harmful content detection and high-confidence fact verification [229, 230].
@@
The low level of generalizability indicates that the agentic AI system is used in specific countries, industries, or even organizations and enterprises. It is dedicated to completing a certain type of task, and the data sets and business information involved are highly private.
Appendix B Datasets for Value Alignment and Evaluation
@@
Table B-II summarizes the existing open-source datasets for value alignment and evaluation. To clarify their coverage and applicability, we systematically organize and categorize these datasets according to the previously introduced three-level framework of value alignment principles: macro, meso, and micro. This hierarchical classification enables a more nuanced understanding of how current resources support value alignment and evaluation at different levels.
@@
Although existing datasets address multiple levels of value alignment, several limitations persist. At the macro level, many datasets—such as Moral Stories [378] and ETHICS [366]—are predominantly grounded in Western cultural contexts, lacking comprehensive representation of alternative value systems, such as those from Asian or African societies. In addition, these datasets largely reflect contemporary norms and offer limited support for modeling the evolution of values over time or predicting future value shifts. At the meso level, datasets such as NaVAB [380], KorNAT [372], and CultureSPA [360] provide valuable insights into national or cultural contexts. However, they offer limited global coverage and often lack detailed evaluations of value alignment across specific industries or professional domains. Micro-level datasets are especially scarce, compared with macro and meso-level resources, they are still markedly underdeveloped, and many remain closed-source. For example, GreedLlama(a financial value-alignment dataset) has not been released to the public for safety reasons [394], and VITAL(an evaluation set for aligning AI with diverse values in healthcare) is likewise unavailable [395]. This situation limits our ability to assess and align value preferences for diverse, fine-grained scenarios in multi-agent systems.
B.1 Dataset Construction Methodologies